# Synthia: Novel Concept Design with Affordance Composition

### v2

@@ Abstract
Text-to-image (T2I) models enable rapid concept design, making them widely used in AI-driven design. While recent studies focus on generating semantic and stylistic variations of given design concepts, functional coherence–the integration of multiple affordances into a single coherent concept–remains largely overlooked. In this paper, we introduce Synthia, a framework for generating novel, functionally coherent designs based on desired affordances. Our approach leverages a hierarchical concept ontology that decomposes concepts into parts and affordances, serving as a crucial building block for functionally coherent design. We also develop a curriculum learning scheme based on our ontology that contrastively fine-tunes T2I models to progressively learn affordance composition while maintaining visual novelty. To elaborate, we (i) gradually increase affordance distance, guiding models from basic concept-affordance association to complex affordance compositions that integrate parts of distinct affordances into a single, coherent form, and (ii) enforce visual novelty by employing contrastive objectives to push learned representations away from existing concepts. Experimental results show that Synthia outperforms state-of-the-art T2I models, demonstrating absolute gains of 25.1% and 14.7% for novelty and functional coherence in human evaluation, respectively. Code is available at https://github.com/HyeonjeongHa/SYNTHIA.
@@
Hyeonjeong Ha1thanks:   Equal contribution., Xiaomeng Jin1∗, Jeonghwan Kim1, Jiateng Liu1, Zhenhailong Wang1, Khanh Duy Nguyen1, Ansel Blume1, Nanyun Peng2, Kai-Wei Chang2, Heng Ji1 1University of Illinois Urbana-Champaign 2University of California Los Angeles {hh38, xjin17, hengji}@cs.illinois.edu
@@ 1 Introduction
Imagine a coffee machine with wheels that brews a morning coffee and delivers it to your bed every morning. This example illustrates a novel concept that is atypical and dissimilar to everyday concepts we regularly encounter in our lives. Novel concept synthesis requires an effective fusion of disparate concepts (e.g., coffee machine, trolley), akin to how humans blend ideas across cognitive domains to generate creative innovations Fauconnier and Turner (2002); Han et al. (2018).
@@
Figure 1: Effect of Affordance Sampling on Novel Concept Generation. Our affordance sampling strategy selects disparate affordance pairs within our ontology, promoting novel functional coherence rather than redundant combinations. Baseline models tend to generate existing concepts for close affordances (Fig 1(a)) but struggle with distant pairs, often introducing multiple objects or omitting functions (Fig 1(b)). In contrast, our models consistently generate functionally coherent novel concepts, achieving higher novelty scores for distant affordance pairs.
@@
Existing studies on conceptual design using T2I models have enabled rapid ideation of novel visual concepts (Cai et al., 2023; Ma et al., 2023; Wang et al., 2024; Lin et al., 2025) by identifying user challenges such as interpreting abstract concepts in language to help visualize a novel design concept (Lin et al., 2025), or using large language models (LLMs) to bootstrap initial ideation in texts (Cai et al., 2023; Zhu and Luo, 2023). However, they often naively feed LLM-generated textual prompts into T2I models, relying on simple key phrases or semantic variations of concept description (Cai et al., 2023; Wang et al., 2024). While existing works show that T2I models can generate images that seem to correctly reflect complex human-formulated textual descriptions (e.g., “beautiful rendering of neon lights in futuristic cyberpunk city”), they do not focus on whether a model can synthesize a novel concept when given a set of affordances (e.g., brew, deliver) as input, while ensuring these affordances are preserved.
@@
An important aspect lacking in existing approaches to concept synthesis is their focus on pixel-based control, overlooking the structural and functional roles embedded in design. Many real-world concepts are naturally “decomposable” into parts, where each part signals a specific functionality. To address this, we propose Synthia, a framework for Concept Synthesis with Affordance composition that generates functionally coherent and visually novel concepts given a set of desired affordances. Unlike prior works relying on complex descriptive text to generate stylistic variations or aesthetic features Richardson et al. (2024); Vinker et al. (2023), Synthia leverages affordances–defined as “the functionality offered by an object or its parts”–as a structural guide for novel concept synthesis. By aligning textual descriptions with affordances as control signals, our models implicitly learn to “decompose and reassemble” functional parts, ensuring that, for instance, a hybrid of a coffee machine and a trolley not only appears novel but also retains its brewing and mobility functions, achieving functional coherence.
Figure 2: SYNTHIA: Novel Concept Design with Affordance Composition. SYNTHIA comprises three stages: (1) Affordance composition curriculum construction, (2) Affordance-based curriculum learning, and (3) Evaluation. In the first stage, we build a training curriculum through sampling affordance pairs from our ontology by gradually increasing the affordance distances. Using our curriculum, we fine-tune T2I models, where they first learn concept-affordance relation from easy data, then integrate multiple affordances into a single functional form from hard data. We employ a contrastive objective with positive (affordances), negative (concepts) constraints, and corresponding images, enforcing visual novelty different from existing concepts. Finally, we evaluate models through automatic evaluation and human evaluation with four metrics: faithfulness, and novelty, practicality, coherence.
@@
To facilitate structured affordance composition, we construct a hierarchical concept ontology that decomposes visual concepts (e.g., Furniture-Sofa) into their constituent parts (e.g., leg, cushion) and associated affordances (e.g., support, rest). It provides a structured representation of concept-affordance associations, serving as the foundation for generating functionally meaningful designs. Inspired by the theory of combinational creativity in humans Han et al. (2018), which suggests novel concepts emerge from disparate ideas, we propose an affordance sampling mechanism that strategically selects affordances associated with sufficiently different concepts using our novel similarity-based metric (§3.1). This ensures that generated designs integrate novel functionalities, avoiding trivial combinations, whereas random sampling yields similar affordances (e.g., cook, heat) that result in redundant outputs (Fig 1(a)).
@@
We also introduce a new curriculum learning scheme that fine-tunes T2I models to progressively learn affordance composition while maintaining visual novelty. Our curriculum gradually increases the affordance distance, allowing models to first learn basic concept-affordance associations from close affordance pairs before tackling complex affordance compositions that integrate multiple affordances into a single, coherent form. To further ensure novelty, we employ contrastive objectives to push learned representations away from existing concepts in our ontology. This addresses a critical limitation of existing T2I models, which struggle to generate coherent multi-functional concepts (Fig. 1(b)). Without structured affordance composition, models tend to default to familiar objects–e.g., when prompted with drive and vacuum affordances, Stable Diffusion models simply generate a car with missing vacuum functions (Fig. 1(b)), rather than blending both affordances into an integrated design. Importantly, unlike existing AI-driven design frameworks that rely on detailed LLM-generated descriptions, Synthia enables direct affordance-based prompting, e.g., “a new design that has functions of {desired affordances}.”. Our model implicitly learns concept-affordance associations, producing novel, structured designs without redundant textual prompting.
@@ 2 Related Work
The advancement of text-to-image (T2I) models has enabled high-quality image synthesis from textual descriptions (Sohn et al., 2023; Xue et al., 2024; Shi et al., 2024; Chen et al., 2024). Especially, the invention of diffusion-based models, such as DALL-E (Ramesh et al., 2021a) and Stable Diffusion (Rombach et al., 2022), significantly increases the performance of the T2I generation by utilizing a transformer-based architecture, where the image embeddings and text encodings are aligned in the shared representation space. For instance, Bao et al. (2024) propose a compositional fine-tuning method for T2I Diffusion Models that focuses on two novel objectives and performs fine-tuning on critical parameters. However, these models still struggle to understand practical functionalities and integrate multiple components into coherent novel concepts. This highlights the need for a new framework to enhance the compositional reasoning ability of T2I models, which our work aims to address.
@@
The great power of T2I models provides a potential boost to content creation (Ko et al., 2023; Rangwani et al., 2024; Sankar and Sen, 2024; Rahman et al., 2024; Tang et al., 2024). Novel concept generation aims to produce visual outputs that extend the existing concepts by specifying the requirements as input to the T2I models. For instance, Concept Weaver (Kwon et al., 2024) first generates a template image based on a text prompt, then refines it using a concept fusion strategy. ConceptLab (Richardson et al., 2024) utilizes Diffusion Prior models and formulates the generation process as an optimization process over the output space of the diffusion prior. Yet, they focus on concept-level generation and ignore the relationships between concepts and their parts. By prioritizing aesthetics, they limit real-world practicality. Our work bridges this gap by designing an affordance-driven framework for novel concept synthesis, ensuring the fusion of desired functions to output novel but practical concepts.
@@ 3 SYNTHIA: Novel Concept Design with Affordance Composition
Our ultimate goal is to utilize T2I models in designing novel concepts that are both visually novel and functionally coherent. Specifically, we take the desired affordances as text inputs, the T2I models should generate an image that depicts the novel concept design. To achieve this, we (1) construct a training recipe that explicitly embeds hierarchical relations on visual concepts, parts, and corresponding affordances, and (2) fine-tune T2I models with curriculum-based optimization.
@@ 3.1 Affordance Composition Curriculum
The primary challenge in the novel concept generation of existing T2I models is the lack of structured functional grounding. These models often struggle to design visually novel yet functionally coherent concepts while maintaining intended functionalities. For example, when combining affordances like Brew and Cut, they may prioritize aesthetics over functionality, omitting parts or objects relevant to Brew (Fig 1(b)). To address this, we construct a structured training recipe in two key steps: (1) building a hierarchical concept ontology, and (2) designing an affordance sampling strategy for curriculum-based training. This improves the model’s composition ability by learning the connection between concepts and affordances.
@@
To provide a structured basis for novel concept synthesis with functional coherence, we define a hierarchical concept ontology that decomposes visual concepts into constituent parts and their affordances, capturing concept-affordance associations (Fig 5). This ontology allows T2I models to retrieve relevant parts based on affordances, enabling generation to be well-grounded on the functionality of concepts rather than superficial visual feature combinations. Formally, we define the ontology as a four-level hierarchy $\mathcal{O}=(\mathcal{S},\mathcal{C},\mathcal{P},\mathcal{A})$. Superordinate $\mathcal{S}$ denotes the highest-level categories, such as furniture, followed by Concept $\mathcal{C}$, which is the set of visual concepts. Each $c\in\mathcal{C}$ belongs to a superordinate category $s\in\mathcal{S}$, e.g., $\mathcal{S}_{\texttt{table}}=\texttt{furniture}$, and decomposes into its parts $\mathcal{P}$ that serve specific functions in an object design, e.g., $\mathcal{P}_{\texttt{table}}=\{\texttt{leg},\texttt{drawer}\}$. The Affordance $\mathcal{A}$ describes functionalities of concepts and parts. Both a concept $c\in C$ and its part $p\in P$ are linked to affordances set $\mathcal{A}_{c}=\{a_{1},\cdots,a_{n}\}\in\mathcal{A}$, e.g, $\mathcal{A}_{\texttt{table}}=\{\texttt{write},\texttt{organize}\}$, and $\mathcal{A}_{p}=\{p_{1},\cdots,p_{n}\}$, e.g., $\mathcal{A}_{\texttt{leg}}=\{\texttt{support}\}$. Our ontology spans 30 superordinates, 590 concepts, 1172 parts, and 686 affordances, explicitly providing a structured representation of how affordance is associated with fine-grained parts for functionally grounded novel concept synthesis.
@@
Given our ontology, we can utilize it to create fine-tuning data to improve the functional coherence of the generated novel concept by T2I models. A naive approach to obtaining training data would be to exhaustively pair all possible affordances. However, this would yield 235K affordance pairs, which is computationally expensive. Moreover, random combination risks generating redundant concepts (e.g., Heat and Cook examples in Fig 1(a)) or functionally incoherent objects. To achieve sufficiently different affordance pairs that enable novel concept synthesis while still being functionally integrable, we introduce a distance-based affordance sampling strategy that selects meaningful, disparate affordance pairs based on ontology-derived distances.
We define a concept distance $D_{C}(c_{i},c_{j})$ between two concepts $c_{i},c_{j}\in\mathcal{C}$ by incorporating functional relatedness at the affordance level and semantic similarity at the concept level. We compute functional relatedness using Jaccard similarity $J(X,Y)=\frac{|X\cap Y|}{|X\cup Y|}$ between their affordance sets while quantifying the semantic similarity Sim via measuring embedding similarity using the BERT (Devlin et al., 2019) model as follows:
$D_{\mathcal{C}}(c_{i},c_{j})$ $=\alpha*\{J(\mathcal{A}_{c_{i}},\mathcal{A}_{c_{j}})+J(\mathcal{A}_{P_{c_{i}}},\mathcal{A}_{P_{c_{j}}})\},$
@@
We further obtain affordance distance $D_{\mathcal{A}}(a_{i},a_{j})$ between two affordances $a_{i},a_{j}\in\mathcal{A}$ by averaging pairwise concept distances $D_{\mathcal{C}}(\cdot,\cdot)$ between associated concepts:
@@
[…]
@@
We sample 600 affordance pairs uniformly across the full distance spectrum and categorize them into three groups. For training images used as pseudo novel concepts, we generate 10 images per pair using DALL-E Ramesh et al. (2021b) with GPT-4 (OpenAI, 2024) generated captions that describe different novel designs integrating the specified affordances. We then filter images using CLIP similarity scores and manually select the top three. This curriculum-based training enables T2I models to learn basic concept-affordance associations while fusing affordances into coherent and functionally meaningful designs. Thus, the T2I models can successfully produce novel concepts that are visually distinctive and functionally coherent.
3.2 Contrastive Fine-tuning with Curriculum Learning

### v4

@@ Abstract
Text-to-image (T2I) models enable rapid concept design, making them widely used in AI-driven design. While recent studies focus on generating semantic and stylistic variations of given design concepts, functional coherence—the integration of multiple affordances into a single coherent concept—remains largely overlooked. In this paper, we introduce Synthia, a framework for generating visually novel and functionally coherent designs based on desired affordances. Our approach leverages a hierarchical concept ontology that decomposes concepts into parts and affordances, serving as a crucial building block for functionally coherent design. We also develop a curriculum learning scheme based on our ontology that contrastively fine-tunes T2I models to progressively learn affordance composition while maintaining visual novelty. To elaborate, we (i) gradually increase affordance distance, guiding models from basic concept-affordance association to complex affordance compositions that integrate parts of distinct affordances into a single, coherent form, and (ii) enforce visual novelty by employing contrastive objectives to push learned representations away from existing concepts. Experimental results show that Synthia outperforms state-of-the-art T2I models, demonstrating absolute gains of 25.1% and 14.7% for novelty and functional coherence in human evaluation, respectively. Code is available at https://github.com/HyeonjeongHa/SYNTHIA.
@@
Hyeonjeong Ha1thanks:   Equal contribution., Xiaomeng Jin1∗, Jeonghwan Kim1, Jiateng Liu1, Zhenhailong Wang1, Khanh Duy Nguyen1, Ansel Blume1, Nanyun Peng2, Kai-Wei Chang2, Heng Ji1 1University of Illinois Urbana-Champaign 2University of California Los Angeles {hh38, xjin17, hengji}@illinois.edu
@@ 1 Introduction
Imagine a coffee machine with wheels that brews your morning coffee and delivers it directly to your bed. This example illustrates a novel concept that is atypical and dissimilar to everyday objects we regularly encounter. Such novel concept design is defined as the synthesis of concepts that are visually unique and functionally coherent through the effective fusion of disparate concepts (e.g., coffee machine, trolley) based on the user’s desired affordances (e.g., brew, deliver). This process mirrors how humans blend ideas across cognitive domains to generate creative innovations fauconnier2002wwt; Han2018THECD.
@@
Figure 1: Effect of Affordance Sampling on Novel Concept Generation. Our affordance sampling strategy selects disparate affordance pairs within our ontology, promoting novel functional coherence rather than redundant combinations. Baseline models tend to generate existing concepts for close affordances (Fig. 1(a)) but struggle with distant pairs, often introducing multiple objects or omitting functions (Fig. 1(b)). In contrast, our models consistently generate functionally coherent novel concepts, achieving higher novelty scores for distant affordance pairs.
@@
Existing studies on conceptual design using T2I models have enabled rapid ideation of novel visual concepts (cai2023designaid; ma2023conceptual; wang2024inspired; lin2025inkspire) by identifying user challenges such as interpreting abstract concepts in language to help visualize a novel design concept (lin2025inkspire), or using large language models (LLMs) to bootstrap initial ideation in texts (cai2023designaid; zhu2023generative). However, they often naively feed LLM-generated textual prompts into T2I models, relying on simple key phrases or semantic variations of concept description (cai2023designaid; wang2024inspired). While existing works show that T2I models can generate images that seem to correctly reflect complex human-formulated textual descriptions (e.g., “beautiful rendering of neon lights in futuristic cyberpunk city”), they do not focus on whether a model can synthesize a novel concept when given a set of affordances as input, while ensuring these affordances are preserved.
@@
An important aspect lacking in existing approaches to concept synthesis is their focus on pixel-based control, overlooking the structural and functional roles embedded in design. Many real-world concepts are naturally “decomposable” into parts, where each part signals a specific functionality. To address this, we propose Synthia, a framework for Concept Synthesis with Affordance composition that generates functionally coherent and visually novel concepts given a set of desired affordances. Unlike prior works relying on complex descriptive text to generate stylistic variations or aesthetic features richardson2024conceptlab; vinker2023concept, Synthia leverages affordances–defined as “the functionality offered by an object or its parts”—as a structural guide for novel concept synthesis. By aligning textual descriptions with affordances as control signals, our models implicitly learn to “decompose and reassemble” functional parts, ensuring that, for instance, a hybrid of a coffee machine and a trolley not only appears novel but also retains its brewing and mobility functions, achieving functional coherence.
@@
To facilitate structured affordance composition, we construct a hierarchical concept ontology that decomposes visual concepts (e.g., Sofa) into their constituent parts (e.g., leg, cushion) and associated affordances (e.g., support, rest). It provides a structured representation of concept-affordance associations, serving as the foundation for generating functionally meaningful designs. Inspired by the theory of combinational creativity in humans Han2018THECD, which suggests novel concepts emerge from disparate ideas, we propose an affordance sampling mechanism that strategically selects affordances associated with sufficiently different concepts using our novel similarity-based metric (§3.1). This ensures that generated designs integrate novel functionalities, avoiding trivial combinations, whereas random sampling yields similar affordances (e.g., cook, heat) that result in redundant outputs (Fig. 1(a)).
@@
We also introduce a new curriculum learning scheme that fine-tunes T2I models to progressively learn affordance composition while maintaining visual novelty. Our curriculum gradually increases the affordance distance, allowing models to first learn basic concept-affordance associations from close affordance pairs before tackling complex affordance compositions that integrate multiple affordances into a single, coherent form. To further ensure novelty, we employ contrastive objectives to push learned representations away from existing concepts in our ontology. This addresses a critical limitation of existing T2I models, which struggle to generate coherent multi-functional concepts (Fig. 1(b)). Without structured affordance composition, models tend to default to familiar objects—e.g., when prompted with drive and vacuum affordances, Stable Diffusion models simply generate a car with missing vacuum functions (Fig. 1(b)), rather than blending both affordances into an integrated design. Importantly, unlike existing AI-driven design frameworks that rely on detailed LLM-generated descriptions, Synthia enables direct affordance-based prompting, e.g., “a new design that has functions of {desired affordances}.”. Our model implicitly learns concept-affordance associations, producing novel, structured designs without redundant textual prompting.
@@ 2 Related Work
The advancement of text-to-image (T2I) models has enabled high-quality image synthesis from textual descriptions (sohn2023styledrop; xue2024raphael; shi2024instantbooth; chen2024pixart; zhou2025contrastive). Especially, the invention of diffusion-based models, such as DALL-E (dalle) and Stable Diffusion (stabled), significantly increases the performance of the T2I generation by utilizing a transformer-based architecture, where the image embeddings and text encodings are aligned in the shared representation space. For instance, bao2024separateandenhancecompositionalfinetuningtext2image propose a compositional fine-tuning method for T2I Diffusion Models that focuses on two novel objectives and performs fine-tuning on critical parameters. However, these models still struggle to understand practical functionalities and integrate multiple components into coherent novel concepts. This highlights the need for a new framework to enhance the compositional reasoning ability of T2I models, which our work aims to address.
@@
The great power of T2I models provides a potential boost to content creation (ko2023large; rangwani2024crafting; sankar2024novelideagenerationtool; rahman2024visual; tang2024lgm; zhou2025contrastivevisualdataaugmentation). Novel concept generation aims to produce visual outputs that extend the existing concepts by specifying the requirements as input to the T2I models. For instance, Concept Weaver (kwon2024concept) first generates a template image based on a text prompt, then refines it using a concept fusion strategy. ConceptLab (richardson2024conceptlab) utilizes Diffusion Prior models and formulates the generation process as an optimization process over the output space of the diffusion prior. Yet, they focus on concept-level generation and ignore the relationships between concepts and their parts. By prioritizing aesthetics, they limit real-world practicality. Our work bridges this gap by designing an affordance-driven novel concept synthesis framework, integrating desired functions to output novel but practical concepts.
@@ 3 Synthia: Novel Concept Design with Affordance Composition
Figure 2: Synthia: Novel Concept Design with Affordance Composition. Synthia comprises three stages: (1) Affordance composition curriculum construction, (2) Affordance-based curriculum learning, and (3) Evaluation. In the first stage, we build a training curriculum through sampling affordance pairs from our ontology by gradually increasing the affordance distances. Using our curriculum, we fine-tune T2I models, where they first learn concept-affordance associations from easy data, then integrate multiple affordances into a single functional form from hard data. We employ a contrastive objective with positive (affordances), negative (concepts) constraints, and corresponding images, enforcing visual novelty different from existing concepts. Finally, we evaluate models through automatic evaluation and human evaluation with four metrics: faithfulness, and novelty, practicality, coherence.
Our ultimate goal is to utilize T2I models in designing novel concepts that are both visually novel and functionally coherent. Specifically, we take the desired affordances as text inputs, the T2I models should generate an image that depicts the novel concept design. To achieve this, we (1) construct a training recipe that explicitly embeds hierarchical relations on visual concepts, parts, and corresponding affordances, and (2) fine-tune T2I models with curriculum-based optimization (Fig. 2).
@@ 3.1 Affordance Composition Curriculum
The primary challenge in the novel concept generation of existing T2I models is the lack of structured functional grounding. These models often struggle to design visually novel yet functionally coherent concepts while maintaining intended functionalities. For example, when combining affordances like Brew and Cut, they may prioritize aesthetics over functionality, omitting parts or objects relevant to Brew (Fig. 1(b)). To address this, we construct a structured training recipe in two key steps: (1) building a hierarchical concept ontology, and (2) designing an affordance sampling strategy for curriculum-based training. This improves the model’s composition ability by learning the connection between concepts and affordances.
@@
To provide a structured basis for novel concept synthesis with functional coherence, we define a hierarchical concept ontology that decomposes visual concepts into constituent parts and their affordances, capturing concept-affordance associations (Fig. 6). This ontology allows T2I models to retrieve relevant parts based on affordances, enabling generation to be well-grounded on the functionality of concepts rather than superficial visual feature combinations. Formally, we define the ontology as a four-level hierarchy $\mathcal{O}=(\mathcal{S},\mathcal{C},\mathcal{P},\mathcal{A})$. The superordinate $\mathcal{S}$ denotes the highest-level categories, such as furniture, followed by Concept $\mathcal{C}$, which is the set of visual concepts. Each $c\in\mathcal{C}$ belongs to a superordinate category $s\in\mathcal{S}$, e.g., $\mathcal{S}_{\texttt{table}}=\texttt{furniture}$, and decomposes into its parts $\mathcal{P}$ that serve specific functions in an object design, e.g., $\mathcal{P}_{\texttt{table}}=\{\texttt{leg},\texttt{drawer}\}$. The Affordance $\mathcal{A}$ describes functionalities of concepts and parts. Both a concept $c\in C$ and its part $p\in P$ are linked to affordances set $\mathcal{A}_{c}=\{a_{1},\cdots,a_{n}\}\in\mathcal{A}$, e.g, $\mathcal{A}_{\texttt{table}}=\{\texttt{write},\texttt{organize}\}$, and $\mathcal{A}_{p}=\{p_{1},\cdots,p_{n}\}$, e.g., $\mathcal{A}_{\texttt{leg}}=\{\texttt{support}\}$. Our ontology spans 30 superordinates, 590 concepts, 1172 parts, and 686 affordances, explicitly providing a structured representation of how affordance is associated with fine-grained parts for functionally grounded novel concept synthesis.
@@
Given our ontology, we can utilize it to create fine-tuning data to improve the functional coherence of the novel concept generated by the T2I models. A naive approach to obtaining training data would be to exhaustively pair all possible affordances. However, this would yield 235K affordance pairs, which is computationally expensive. Moreover, random combination risks generating redundant concepts (e.g., Heat and Cook examples in Fig. 1(a)) or functionally incoherent objects. To achieve sufficiently different affordance pairs that enable novel concept synthesis while still being functionally integrable, we introduce a distance-based affordance sampling strategy that selects meaningful, disparate affordance pairs based on ontology-derived distances.
We define a concept distance $D_{C}(c_{i},c_{j})$ between two concepts $c_{i},c_{j}\in\mathcal{C}$ by incorporating functional relatedness at the affordance level and semantic similarity at the concept level. We compute functional relatedness using Jaccard similarity $J(X,Y)=|X\cap Y|/|X\cup Y|$ between their affordance sets while quantifying semantic similarity Sim by measuring embedding similarity using the BERT (devlin2019bertpretrainingdeepbidirectional) model as follows:
$D_{\mathcal{C}}(c_{i},c_{j})$ $=\alpha*\{J(\mathcal{A}_{c_{i}},\mathcal{A}_{c_{j}})+J(\mathcal{A}_{P_{c_{i}}},\mathcal{A}_{P_{c_{j}}})\}$
@@
We further obtain the affordance distance $D_{\mathcal{A}}(a_{i},a_{j})$ between two affordances $a_{i},a_{j}\in\mathcal{A}$ by averaging the pairwise concept distances $D_{\mathcal{C}}(\cdot,\cdot)$ between associated concepts:
@@
Based on our distance metric, close affordance pairs associated with similar concepts, e.g., $\{\texttt{sit},\texttt{rest}\}$ from $\{\texttt{sofa},\texttt{chair}\}$, support learning basic affordance-concept associations, which can be easily merged into existing concepts. In contrast, distant affordance pairs derived from sufficiently distant concepts, e.g., $\{\texttt{drive},\texttt{vacuum}\}$ from $\{\texttt{car},\texttt{vacuum cleaner}\}$, enforce greater functional coherence by requiring meaningful part-affordance integration, which is more complex than a trivial combination.
@@
We sample 600 affordance pairs uniformly across the full distance spectrum and categorize them into three groups. For training images used as pseudo novel concepts, we generate 10 images per pair using DALL-E ramesh2021zeroshottexttoimagegeneration with GPT-4o (openai2024gpt4ocard) generated captions that describe different novel designs integrating the specified affordances (Details are described in the Appendix B.1). We then filter images using CLIP similarity scores and manually select the top three. This curriculum-based training enables T2I models to learn basic concept-affordance associations while fusing affordances into coherent and functionally meaningful designs. Thus, the T2I models can successfully produce novel concepts that are visually distinctive and functionally coherent.
3.2 Affordance-based Curriculum Learning